Question 11 Normal mu 2.09 sigma 0.21 xi P(X<=xi) 1.61 0.0111 1.76 0.0580 2.39 0.9234 2.53 0.9819 P(X<=xi) xi 0.10 1.8209 0.20 1.9133 0.30 1.9799 0.40 2.0368 Normal mu 2.09 sigma 0.24 xi P(X<=xi) 1.61 0.0228 1.76 0.0846 2.39 0.8944 2.53 0.9666 P(X<=xi) xi 0.10 1.7824 0.20 1.8880 0.30 1.9641 0.40 2.0292 The mean weight for a part made using a new production process is 2.09 pounds. Assume that a normal distribution applies and that the standard deviation is 0.24 pounds. Based on this paragraph of text, use the correct excel output above to answer the following question. Find the probability that weight for a part made using the new production process is less than 1.76 or greater than 2.39 pounds. a. 0.8098 b. 0.8654 c. 0.1902 d. None of the answers is correct e. 0.1346
Answer #1
for normal distribution z score =(X-μ)/σx
here mean=       μ= 2.09
std deviation   =σ= 0.240

P(X<1.76)+P(X>2.39) =0.08456+(1-0.8944)=0.1902
option C is correct